OpenStudy (anonymous):

can anyone simplify 4m^3/3m^2n X(times) 9n^2/30m^2n^2

jhonyy9 (jhonyy9):

\[4m ^{3}/3m ^{2}n * 9n ^{2}/30m ^{2}n ^{2}=4m/3n *3/10m ^{2}=\]

jhonyy9 (jhonyy9):

=2/n * 1/5m = 2/(5mn)